I'm trying to help my 8th grade daughter study for math final in algebra I and I did bad on word problems when I was a freshman so I need help helping her please. The question is this:
4 children have a total of 92 marbles
Bill has 8 more than Allen
Charles has 2 times the number of Bill
Daryl has 3 times the number of Charles
How many marbles does Charles have?
She got as far as this:
8 + A + A + 2B + 3C = 92
8 + 2A + 2B + 3C = 92
Then we are stuck

Answer by DrBeeee(684)

You have that right idea, but stalled out after Bill.
Let A = Allen's marbles.
Then
(1) B = A + 8 and
(2) C = 2*B or
(3) C = 2*(A + 8) or
(4) C = 2*A + 16 and
(5) D = 3*C or
(6) D = 3*(2*A + 16) or
(7) D = 6*A + 48
Now add
(8) A + (A + 8) + (2*A + 16) + (6*A + 48) = 92 or
(9) 10*A = 92 -8 -16 -48 or
(10) 10*A = 92 - 72 or
(11) A = 2
Then
(12) B = 10
(13) C = 20
and
(14) D = 60
Is(2 + 10 + 20 + 60 = 92)?
Is (92 = 92)? Yes
Answer: Allen has 2 marbles, Bill has 10, Charles has 20 and Daryl has 60.